OpenStudy (younginslim):

Write the equation of a line with a slope of 4 and a y-intercept of −3. 4x + y = −3 4x – y = −3 y = −3x + 4 y = 4x − 3

OpenStudy (reemii):

The slope is the "rate of increase".  That means: for \(x=0\), the corresponding value of \(y\) is the intercept. Which of your equations verifies : "if x=0, then y=-3" ? (two possibilities)

OpenStudy (younginslim):

4x + y = -3

OpenStudy (reemii):

good, there's another one: y = 4x - 3.

OpenStudy (younginslim):

So it's D?

OpenStudy (reemii):

That's two candidates. Among those two, which one has a slope equal to 4 ? To find the slope based on an equation, always isolate the \(y\), then read the factor of \(x\). Example: \(3x+2y=1\) becomes \(y = \frac32x + \frac12\). Factor of \{x\) is \(\frac32\), it is the slope. What about the two candidates?

OpenStudy (reemii):

Oh well, yes

OpenStudy (younginslim):

Ohhh I see

OpenStudy (reemii):

(oops. I meant: \(y = -\frac32x + 1\)... )

OpenStudy (younginslim):

So the answer is still the same?

OpenStudy (reemii):

that was just my example.. \(3x+2y=1\) us equivalnet to \(y=-\frac32x+1\) and the slope is \(-\frac32\). For your question, -slope of line with equation 4x + y = −3 is \(-4\). -slope of line with equation y = 4x-3 is .. \(4\) (nothing to do here). -> answer D)

OpenStudy (reemii):

(aaargh \(y = -\frac32x + \frac12\) but that doesn't matter!! ;))

OpenStudy (younginslim):

Oh thanks. Can u help me with 2 more?

OpenStudy (reemii):

2 like this one sure

OpenStudy (reemii):

and I'm not alone :'>

OpenStudy (younginslim):

What you mean?

OpenStudy (reemii):

There are other people who are able to answer. Please post your question

OpenStudy (younginslim):

Ok

OpenStudy (skullpatrol):

Close this question and post a new one please :-)

OpenStudy (younginslim):

ok

OpenStudy (skullpatrol):

Thanks